OpenStudy (anonymous):

Solve the following system of equations by using the substitution method. y = x + 3 y = 2x – 4

jimthompson5910 (jim_thompson5910):

y = x + 3 ... Start with the first equation 2x - 4 = x + 3 ... replace y with 2x - 4 (since y = 2x-4) 2x - 4 - x = 3 ... Subtract x from both sides. 2x - x = 3 + 4 ... Add 4 to both sides. 1x = 7 ... Combine like terms. x = 7 ... use the idea that 1x is the same as x So x = 7 Now use this value of x to get the value of y. I'll let you do this part.

OpenStudy (anonymous):

value of y is 10?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome